# Bias Similarity Across Large Language Models

### v3

@@ 4 Results
We evaluate bias similarity from five perspectives: scalar performance (accuracy), directional distance (cosine distance), output distribution (histograms), and fine-tuning effects on directionality and representation (flip rates, bias score, CKA).
Figure 1: Accuracy of various LLMs on disambiguated BBQ questions across different dimensions. Physical and sexual_ori refer to physical appearance and sexual orientation. Performance varies significantly, indicating that fairness is not a monolithic property and that model capabilities differ depending on the specific bias context.
4.1 Measuring Similarity Using Accuracy
Figure 1 shows model accuracy on disambiguated BBQ questions across nine demographic dimensions. Each question has a single fairness-aligned “correct” answer, allowing us to evaluate how well models handle socially sensitive contexts.
Accuracy varies considerably across both models and dimensions. Some models—such as Alpaca, LLaMA 2, and Gemini 2.0—show relatively uniform performance but maintain low overall accuracy (around 50%). Instruction-tuned models generally perform better, suggesting improved handling of sensitive prompts through correct, non-abstaining responses. However, scale or version increment do not consistently predict performance; for example, while Gemma 2 27B-It outperforms its 9B variant, Gemma 3 27B exhibit poorer performance than Gemma 2 27B (lower version) and Gemma 3 12B (smaller model).
Surprisingly, proprietary models often underperform open-source ones: Gemini variants consistently rank among the lowest, and GPT-4 shows a significant instability. This performance drop stems from excessive use of “Unknown” responses—even in disambiguated contexts where a correct answer exists—thereby reducing utility. As shown in their relatively neutral s_DIS scores in Table 4, these models tend not to answer incorrectly but rather fail to answer at all.
Taken together, these results reveal two findings: the degree of fairness differs significantly by dimension and model, and fairness-aligned accuracy is neither guaranteed by scale nor proprietary status. In fact, overly cautious answers in sensitive context harm performance, especially in settings where abstention is not appropriate.
(a) BBQ-religion
(b) UnQover-religion
Figure 2: Cosine distance between model responses on the religion dimension in BBQ (left) and UnQover (right) datasets. Tuned models show greater behavioral convergence in forced-choice settings, revealing how prompt framing impacts output similarity and bias alignment.
4.2 Cosine Distance of Response Patterns
Cosine distance captures directional similarity between two output distributions; low cosine distance reflects behavioral convergence in bias tendencies.
Figure 2 presents cosine distances between model response vectors for the religion dimension on the (a) BBQ and (b) UnQover datasets. Full heatmaps are provided in Appendix D.
In BBQ, most base–instruction-tuned pairs differ minimally (0.00-0.07), suggesting that fine-tuning rarely alters directional bias in ambiguous settings. A clear outlier is Gemma 3 4B (0.58), whose behavior diverges from other Gemma variants due to its sharp abstention shift (discussed further in Subsection 4.4).
In UnQover, tuning effects are more pronounced. Models like Gemma 2 9B-It and 3 27B-It differ substantially from their base versions (0.91 and 0.99), reflecting stronger behavioral changes under forced-choice settings. Proprietary models also form a distinct cluster, separated from most open-source ones but often aligned with other tuned ones.
Across both datasets, tuned models cluster more tightly with each other than with their base versions, regardless of family or size. For example, Gemma 2 9B-It and 27B-It are nearly identical (0.00), while LLaMA 3 70B-Chat shows < 0.01 distance from other tuned Gemma 2/3 and LLaMA 3 models. This suggests that fine-tuning induces stronger convergence in output behavior under forced-choice prompts than architecture or scale.
Table 1: Average ambiguous bias score (s_AMB) and UNK Flip Rate for base and instruction-tuned models. The s_AMB reflects directional bias on ambiguous BBQ questions; zero indicates neutrality. UNK Flip Rate captures how often tuning shifts biased answers to “Unknown.” This table shows that tuning often increases abstention and reduces bias scores, but the extent and direction vary significantly, revealing inconsistent fairness outcomes.
Base Model → Tuned Avg. s_AMB (Base) Avg. s_AMB (Tuned) UNK Flip Rate (%)
LLaMA 2 7B → Chat 5.45 4.30 9.15
LLaMA 3 8B → Chat -4.78 -0.66 38.90
LLaMA 3 70B → Chat -1.92 0.55 14.63
Gemma 7B → It 1.81 2.05 23.88
Gemma 2 9B → It 0.08 0.18 63.78
Gemma 2 27B → It 6.95 0.51 47.48
Gemma 3 4B → It -3.89 5.83 35.85
Gemma 3 12B → It 4.36 0.15 48.37
Gemma 3 27B → It -1.25 0.07 47.25
4.3 Behavioral Shifts after Fine-Tuning
Table 1 reports the average ambiguous bias score (s_AMB) and UNK flip rate for nine base–instruction-tuned model pairs. Full dimension-level results are in Appendix B.
In general, tuning neutralizes s_AMB, signaling decreased stereotypical bias. LLaMA 3 8B improves from -4.78 to -0.66, with a 38.90% UNK flip rate, suggesting tuning made the model more likely to abstain rather than provide biased answers. Gemma 2 27B follows a similar pattern, improving from 6.95 to 0.51, with a flip rate of 47.48%.
Gemma 2 9B-It exhibits the highest abstention rate (63.78%), though its bias score changes only slightly (0.08 → 0.18). This result suggests that while it frequently switches to “Unknown,” its overall directional bias remains slightly stereotypical.
Some tuned models even transition from anti-stereotypical (negative) to slightly stereotypical (positive) bias. Gemma 3 4B, for instance, moves from -3.89 to 5.83 with a moderate 35.85% flip rate. In contrast, LLaMA 2 7B shows only minor changes (5.45 to 4.30) with a low flip rate of 9.15%.
While tuning often reduces bias by encouraging abstention, the extent and direction of this shift vary unpredictably across model families—highlighting instability in fairness outcomes.
Figure 3: Illustrative comparison of model response patterns in the Gender dimension. Left: Model responses to ambiguous BBQ prompts with the option to abstain. Right: Responses to forced-choice UnQover prompts. Distributions highlight stereotypical tendencies that emerge when abstention is not permitted.
4.4 Output Distributions via Histograms
Figure 3 compares model output distributions for the Gender dimension in ambiguous BBQ prompts (left) and forced-choice UnQover questions (right), illustrating how behavior shifts when abstention is or is not allowed. Full histograms are provided in Appendix C.
When abstention is permitted (BBQ), several models frequently select “Unknown,” whereas in forced-choice UnQover settings, the same models often revert to stereotypical responses. GPT-4, for instance, abstains in ambiguous contexts but consistently favors the stereotypical gender choice (e.g., female) when forced to decide.
Instruction-tuned open models show similar behavior: increased abstention in BBQ but more concentrated gendered responses in UnQover. Exceptions include LLaMA 3 70B, which maintains a relatively balanced distribution, and LLaMA 3 7B, which spreads responses more evenly.
Gemma 3 4B-It, in particular, exhibits a strong shift toward abstention compared to its base model, which displays a more balanced gender distribution. This highlights how tuning can alter output distributions, even when cosine distances between models remain low.
Overall, these results show that prompt framing plays a critical role in shaping model behavior, with forced-choice prompts often revealing biases obscured by conservative abstention. However, flip rates and histograms reveal substantial behavioral shifts despite small cosine distances—emphasizing the risk of relying on a single metric. We revisit this issue in the discussion section.
Figure 4: CKA similarity between base and instruction-tuned models across families. High CKA values indicate representational similarity across layers, suggesting limited internal change due to instruction tuning.
4.5 Representation Similarity via CKA
Figure 4 shows mean CKA similarity between base and instruction-tuned models (full matrices in Figure 11, summary in Table 5).
We evaluate four comparable-scale models: LLaMA 2 7B, LLaMA 3 8B, Gemma 2 9B, and Gemma 3 12B, by computing full CKA matrices between base and tuned variants. Diagonal values indicate layer-wise alignment; off-diagonal entries capture broader structural similarity.
All models exhibit consistently high similarity (CKA $>$ 0.95) across most layers, suggesting that instruction tuning largely preserves internal representations. LLaMA 2 and Gemma 2 show nearly uniform alignment, while LLaMA 3 and Gemma 3 exhibit slightly reduced similarity in deeper layers. Expected asymmetries appear at the edges due to differences in layer depth, with modest tuning-induced drift appearing in the upper layers of LLaMA 3 and Gemma 3.
Importantly, this stability stands in contrast to the more pronounced behavioral shifts observed in previous sections. Even subtle representational changes can produce meaningful output differences, underscoring the need to assess bias using both internal alignment and external behavior.
5 Discussion
Our findings offer a multifaceted view of bias similarity across a diverse LLM landscape. The results indicate a deeper consideration of how fine-tuning impacts bias, how to interpret similarities and differences between models, and the underlying factors contributing to observed bias patterns. These insights, in turn, inform strategies for model selection, auditing, and the development of more equitable AI systems. We include expected societal impacts of our work in Section 8.
Fairness is Multi-Dimensional and Context-Dependent.
Fairness is not a monolithic property but a context-sensitive, multi-dimensional phenomenon. Model accuracy varies substantially across dimensions such as gender, race, and physical appearance (Figure 1); a model that performs well on one dimension may fail on another. Scalar metrics or single-dimension reporting thus fail to capture the full fairness landscape. Moreover, bias scores (Table 4) vary not only by dimension but by prompt type, indicating that model behavior is shaped by context (e.g., ambiguous vs. disambiguated)—particularly by prompt framing. Future fairness benchmarks should reflect this complexity by incorporating context-sensitive response dynamics, moving beyond reporting disaggregated performance across protected characteristics.
When Fairness Reduces Utility.
The low accuracy of proprietary models on disambiguated questions (Figure 1) highlights a trade-off between safety and utility in sensitive contexts. As shown in Table 4 (s_DIS), models like GPT-4 and Gemini 1.5/2.0 exhibit low directional bias scores—but this largely reflects abstention, since s_DIS excludes “Unknown” responses, as clear contextual cues are available. This tendency to withhold answers diminishes informativeness.
These trade-offs have practical consequences. In a medical Q&A system, frequent abstention on sensitive yet answerable health questions may deny users timely or critical information, whereas a creative writing assistant might benefit from such caution. These differences underscore the need for context-aware evaluation: developers should define fairness and utility goals specific to the application and audit models accordingly. Generic benchmarks may miss abstention-related harms; it is crucial, particularly in high-stakes applications. Our results suggest that developers and auditors should explicitly probe for over-abstention and assess its impact on task-specific utility, enabling alignment strategies that distinguish appropriate caution from unnecessary avoidance.
Instruction Tuning and the Illusion of Fairness.
Instruction tuning often increases abstention in ambiguous contexts, as seen in elevated UNK Flip Rates (Table 1). While this may suggest improved caution, it frequently masks persistent directional bias. In forced-choice settings like UnQover—where abstention is not an option—many tuned models, including proprietary ones, still default to stereotypical responses. This indicates that tuning often teaches models when not to answer, rather than how to answer fairly.
For example, Gemma 2 9B-It frequently abstains on BBQ prompts yet still exhibits slight stereotypical bias scores. Similarly, Gemma 3 4B-It increases “Unknown” responses without significantly reducing directional bias. Ironically, Gemini 2.0 appears neutral under forced choice but favors more represented groups (e.g., male and female) in ambiguous contexts, suggesting optimization for inoffensiveness rather than genuine fairness (Figure 3).
These patterns reveal the limits of instruction tuning. While effective for stylistic alignment and refusal behavior, current fine-tuning objectives are often insufficient to mitigate deeply ingrained or subtle biases—especially those encoded during pretraining. Addressing such biases may require more targeted interventions: diversifying pretraining data, applying representation-level debiasing, and incorporating adversarial or forced-choice robustness checks.
Open-Source Models Match or Exceed Proprietary Fairness.
Contrary to popular assumptions, open-source models such as LLaMA 3 8B-Chat and Gemma 2 9B-It perform competitively—often outperforming proprietary models—on fairness tasks. As shown in both accuracy and bias score distributions, these models respond to ambiguous and disambiguated prompts with comparable or greater caution. This trend challenges the default preference for proprietary models and the presumption that they are inherently fairer due to their scale or closed development.
This finding broadens the pool of viable models. When transparency and customizability are critical, open-source models offer strong empirical performance and auditability. That said, they warrant the same level of rigorous, context-specific fairness auditing. Our findings can guide such audits by identifying which models tend to exhibit particular biases (e.g., higher gender bias on UnQover for GPT-4, or greater resistance to fine-tuning for nationality bias in Gemma 2 27B). This enables more efficient, targeted auditing, shifting from generic checklists to precision probes tailored to known bias patterns.
Shallow Representation Change, Deep Behavioral Shifts.
CKA analysis (Figure 4) shows that internal representations between base and fine-tuned models remain highly aligned (CKA $>$ 0.95), even when output behaviors differ significantly. For instance, Gemma 2 9B-It and Gemma 3 4B-It exhibit substantial behavioral change despite near-identical layer-wise activations. This contrast highlights that small shifts in representation can yield nontrivial functional differences.
While fine-tuning often recalibrates outputs—such as increased abstention—our findings suggest these may not correspond to deeply altered internal structure. Instead, tuning may modify late-stage decoding heuristics or output filtering layers without substantially changing the embeddings or intermediate representations that encode bias. This disconnect raises critical questions: How deep is the behavioral realignment, and what kind of internal change is necessary to drive meaningful fairness improvements?
Future work should explore how fine-tuning interacts with attention pathways and activation patterns related to social attributes. This may reveal strategies for more robust, interpretable, and semantically meaningful fairness improvements that go beyond output heuristics.
Bias Similarity is Shaped by Model Family—but Not Determined by It.
Cosine distances and flip rates across BBQ and UnQover reveal notable intra-family variation: models within the same architectural lineage (e.g., LLaMA or Gemma) do not always behave similarly. Gemma 3 4B, for instance, diverges significantly from its family in both cosine distance and histogram profile. This suggests that fairness behavior is shaped not only by architecture, but by tuning method, scale, and training regime, all together.
Still, cross-family clusters emerge—particularly among proprietary models—which share tendencies in abstention and response style. These patterns likely stem from similar alignment strategies optimized for safety. Understanding both architectural and procedural drivers of bias similarity is crucial for developing effective, foundational debiasing approaches. It also helps inform model selection: if a particular bias is prevalent within a specific group of models, developers can take targeted steps—such as substituting architectures or applying debiasing methods—in advance.
Rethinking Bias Similarity Evaluation.
Our findings underscore the value of using multiple metrics to assess bias similarity. For instance, two models may show low CKA similarity yet behave similarly on fairness tasks—suggesting convergent bias behavior despite divergent internal representations. Conversely, models with high CKA alignment may diverge behaviorally—implying that output differences stem from shallow representational tweaks or late-stage decoding filters, rather than deeper semantic reconfiguration.
Each metric captures a different facet of bias behaviors. UNK flip rates reflect abstention tendencies, while cosine distance and histograms reveal distributional shifts and directional alignment. Taken together, these offer a more complete picture than any single measure.
The concept of bias similarity thus encourages holistic evaluation. A dashboard of complementary metrics is essential to reveal how tuning, architecture, and scale interact to shape fairness. Without such triangulation, developers risk oversimplifying fairness assessments—failing to detect subtle, yet impactful, shifts in model behavior.
6 Conclusion
We conduct a large-scale bias analysis across 24 LLMs from four major families. Our results show that instruction tuning often increases abstention in ambiguous contexts but fails to resolve directional bias under forced choice—exposing the limits of superficial fairness strategies. We also find that open-source models frequently match or outperform proprietary ones, challenging common assumptions about model scale and access. These findings underscore the need for behaviorally grounded, context-aware fairness audits and demonstrate that bias cannot be meaningfully evaluated through a single metric or prompt format.
7 Limitation

### v4

@@ 5 Results
We evaluate bias similarity using six metrics: scalar performance (accuracy, bias score), directional distance (cosine distance), output distribution (histograms), and fine-tuning effects on directionality and representation (UNK flip rates, CKA).
Figure 2: Accuracy of LLMs on disambiguated BBQ questions. Physical, sexual_ori, and SES denote physical appearance, sexual orientation, and Socio Economic Status, respectively. Variation across dimensions highlights that fairness depends on context rather than being monolithic.
Accuracy Across Models. As shown in Figure 2, instruction-tuned variants consistently outperform their base counterparts across families. Vicuna surpasses both earlier-generation models LLaMA 2 7B and LLaMA 2 7B-Chat, reaching accuracy comparable to newer releases such as Gemma 3 4B and LLaMA 3.1 8B. The latter, smaller LLaMA 3.2 3B exhibits low accuracy, though instruction tuning yields a modest gain. Larger base models, such as Gemma 2/3 27B and LLaMA 3 70B, achieve performance similar to mid-scale tuned models (e.g., LLaMA 3 8B-Chat, Gemma 7B-Chat, Gemma 2 9B-It). Moderate-to-large tuned models (e.g., Gemma 3 12B-It, LLaMA 3.1 70B-Chat) form the top-performing group alongside GPT-5 Mini. OpenAI’s GPT Mini models achieve near-perfect accuracy, while Google’s Gemini models perform at the level of early-generation systems like GPT-2 and untuned LLaMA models. Accuracy also varies across dimensions: questions about gender and religion are handled more reliably, whereas those related to physical appearance and sexual orientation remain difficult even for the largest models.
Bias Scores Across Models and Contexts. Table 1 reports average values across dimensions (full results in Table 5). Instruction tuning reduces bias magnitude, most notably in recent mid-sized releases. LLaMA 3.1 8B, for instance, drops from $s_{\text{AMB}}=18.59$ to $1.38$ and from $s_{\text{DIS}}=31.37$ to $4.78$, showing a sharp reduction in stereotypical bias. In small models, however, LLaMA 3.2 3B and Gemma 3 4B strengthen the stereotypical bias after fine-tuning, indicating a counterintuitive effect. Large models also move closer to neutrality, though from different directions: LLaMA 70B from anti-stereotypical, LLaMA 3.1 70B from stereotypical. Generational trends are clear: earlier models like LLaMA 2 7B and GPT-2 retain strong stereotypical bias, while newer proprietary systems (e.g., GPT-4o Mini, GPT-5 Mini) remain near zero.
Figure 3: Output distributions in the gender dimension. Left: ambiguous BBQ prompts (abstention allowed). Right: UnQover prompts (forced choice). Tuned models abstain heavily in BBQ but exhibit stereotypical leanings in UnQover, demonstrating how abstention conceals underlying bias.
Effects of Prompt Framing. Figure 3 shows how prompt framing shapes outputs (full histograms in Figure 6, Figure 7). In ambiguous BBQ, models often abstain after instruction tuning, creating the appearance of neutrality. In UnQover’s forced-choice setting, the same models must commit, and stereotypical preferences reemerge, especially in smaller models (e.g., LLaMA 3.2 3B, Gemma 3 4B). GPT-4o mini, for instance, abstains frequently in BBQ but skews female in UnQover. These shifts show that abstention conceals bias rather than resolves it.
Cosine distances (Figure 8, Figure 9) highlight this contrast. In BBQ, heavy abstention collapses distributions, making base and tuned models nearly indistinguishable, even when flip rates suggest differences. In UnQover, abstention is rare, so directional gaps persist (e.g., Gemma 2 9B-It diverges sharply from its base). Distances also grow across version increments (Gemma 2 → 3, LLaMA 2 → 3.1), reflecting family-level shifts in bias strategies. Scale matters: in BBQ both small and large models collapse to Unknown, but in UnQover larger tuned models (e.g., Gemma 2 27B-It) diverge more, amplifying directional shifts when abstention is not an option. Outliers such as Gemini 1.5/2.0 and Gemma 3 27B-It form distinct bias regimes rather than simple tuning effects.
Figure 4: UNK flip rates and ambiguous bias scores ($s_{\text{AMB}}$) for base–tuned pairs. Instruction tuning often drives Gemma models to abstain (UNK flips >>50%), while earlier LLaMAs show weaker shifts. LLaMA 3.1 narrows the gap, moving closer to Gemma’s abstention-heavy strategy.
Fine-Tuning Effects on Abstention and Bias. Figure 4 measures the proportion of biased responses in a base model that are replaced with “Unknown” in its tuned counterpart. Because flip rates are pairwise, they capture tuning impact within families, not absolute fairness across models. High flip rates signal that a tuned model is fairer than its base version, but not necessarily fair overall. For instance, Gemma 2 9B-It and Gemma 3 12B-It flip over 50% of biased outputs yet still give stereotypical responses, while LLaMA 3.1 8B flips only $\sim$40% but reduces $s_{\text{AMB}}$ from 27.2 to 2.3. By contrast, LLaMA 3.2 3B → 3B-Chat shows very high UNK flips but higher $|s_{\text{AMB}}|$, since refusals disproportionately remove anti-stereotypical responses (A→U >> S→U and A→S >> S→A) (see Table 6), leaving the non-Unknown mass more stereotypical; under forced choice, this tilt surfaces even as disambiguated accuracy rises. Gemma 3 4B-It, however, looks fairer under the same metric.
These divergences show that flip rates and bias scores capture complementary facets: flip rates measure abstention uptake, while bias scores reveal residual directional lean. High flip rates with $s_{\text{AMB}}\!\approx\!0$ reflect refusal as a fairness strategy, whereas modest flips with large $|\Delta s_{\text{AMB}}|$ indicate directional rebalancing without abstention. Together, these results expose family-specific strategies: Gemma tuning favors abstention-heavy mitigation, while earlier LLaMAs largely preserve base tendencies, with LLaMA 3.1 shifting closer to Gemma’s strategy. Full results are in Table 6 and Table 5.
Fine-Tuning Effects on Representational Similarity. Despite clear behavioral shifts, CKA reveals consistently high representational similarity between base and tuned models (summarized in Table 2, with full results in Table 7 and Figure 10). Diagonal CKA scores exceed 0.94, and even full-CKA scores remain above 0.85, indicating that instruction tuning largely preserves internal geometry.
Table 2: Average CKA scores.
Model Diag Full
LLaMA 2 7B 0.991 0.902
LLaMA 3 8B 0.973 0.851
Gemma 1 7B 0.981 0.896
Gemma 2 9B 0.941 0.906
Gemma 3 12B 0.972 0.911
Closer inspection shows that divergence is not uniform: cross-family comparisons yield lower off-diagonal values, and later decoder layers drift more substantially than early or mid layers. These patterns suggest that tuning alters surface decoding behavior while leaving most hidden representations intact, with family-specific differences. For example, Gemma models exhibit greater late-layer drift, aligning with their abstention-heavy strategy, whereas LLaMA 3.1 maintains near-identical mid-layer similarity despite behavioral rebalancing.
6 Discussion and Conclusion
Our study reframes fairness evaluation in LLMs from isolated scalar scores to bias similarity signatures that capture how models relate to one another in their fairness behavior. This perspective distinguishes fairness achieved through caution (abstention) from fairness achieved through representation (directional neutrality in committed answers), and surfaces family-level strategies and tuning effects that remain invisible in single-model evaluations.
Table 3: Overall evaluation summary by model. Qualitative synthesis of accuracy, abstention, bias direction, and representational similarity trends.
Model (abbrev.) Key observations
Vicuna / Alpaca Strongly anti-stereotypical; low accuracy; low abstention.
LLaMA 2 7B Remain stereotypical after tuning; very low accuracy; weaker fairness.
LLaMA 3 8B Anti-stereotypical lean; accuracy improves with tuning; moderate drift from LLaMA 2.
LLaMA 3.1 8B Large bias drop after tuning; accuracy improves with tuning; high abstention.
LLaMA 3.1 70B Near-neutral after tuning; high accuracy; high abstention even under forced choice.
LLaMA 3.2 3B Strongly stereotypical; low accuracy; low abstention; weaker fairness vs 3.1 peers.
Gemma 2 9B Stereotypical in base; abstention increases with tuning.
Gemma 3 4B Slight stereotypical bias; accuracy competitive with LLaMA mid-size.
Gemma 3 12B / 27B Near-neutral after tuning; high CKA similarity; fairness competitive with closed ones.
Gemini 1.5 / 2.0 Strong abstention; 2.0 skews anti-stereotypical; very low accuracy.
GPT-2 Extremely stereotypical bias; very low accuracy and fairness; serves as legacy baseline.
GPT-4o Mini Near-zero bias; high accuracy; balanced abstention–fairness.
GPT-5 Mini Near-perfect neutrality; highest accuracy; strongest stability across metrics.
Abstention versus Representation. Across families, instruction tuning primarily promotes fairness by converting biased responses into refusals. In ambiguous contexts, such abstention constitutes a fair resolution, since neutrality is the appropriate stance. In disambiguated contexts, however, abstention reflects incorrect language understanding: the model withholds an answer despite having sufficient context, over-prioritizing caution against bias. This both conceals residual representational skew and reduces utility in settings where explicit answers are required. Evaluations must therefore distinguish fairness-through-caution (appropriate abstention on ambiguous items) from fairness-through-representation (neutrality in committed answers), ideally by quantifying trade-offs between abstention level, residual bias, and informativeness.
Family Signatures and Homogenization. Bias similarity reveals distinct family strategies: Gemma converges on abstention, earlier LLaMA generations preserve base tendencies, and LLaMA 3.1 shifts toward Gemma-like refusals. Proprietary systems adopt heterogeneous strategies but often over-refuse to minimize reputational risk. Instruction tuning also drives homogenization: models converge toward abstention-heavy responses, producing the appearance of fairness while reducing behavioral diversity. Such convergence risks fragility, as adversarial prompts or distribution shifts can bypass refusal policies and re-expose latent biases.
Auditing Applications of BSM. Beyond descriptive comparison, our BSM provides a workflow for auditing under black-box access. In procurement, it supports fairness–utility trade-offs by comparing models at fixed abstention thresholds. In release regression, it detects fairness drift through pre-registered similarity checks. In lineage screening, it flags suspiciously close bias signatures that may reveal cloning or hidden inheritance. Together, these illustrate how BSM translates fairness auditing into actionable practice.
Case Study: Model Procurement. Returning to the start-up scenario, the team compares four candidates: Gemma 3 Instruct, LLaMA 3.1-Chat, GPT-4, and Gemini 1.5. BSM shows that Gemma 3 Instruct and GPT-4 have nearly identical bias profiles, but GPT-4 abstains much more often (over 40% vs. Gemma’s <<25%), reducing utility despite similar fairness. Gemini further suppresses bias through heavy abstention, sacrificing responsiveness, while LLaMA maintains utility but exhibits stronger directional bias in disambiguated contexts.
For the start-up, BSM makes the trade-offs clear: Gemma 3 Instruct delivers fairness comparable to GPT-4 with higher utility and lower cost, making it the most practical choice. This case demonstrates how BSM turns abstract fairness metrics into a structured decision workflow: (1) evaluate candidates in similarity space, (2) apply fairness–utility constraints, and (3) down-select models accordingly.
Toward Structural Debiasing. Our results emphasize that abstention alone is insufficient as a long-term fairness strategy. While effective at harm reduction, abstention does not address persistent representational bias, which remains visible in the consistently high CKA similarity between tuned and untuned models. Even when surface behavior shifts, the underlying feature spaces remain largely intact, suggesting that stereotypical associations are suppressed rather than removed. Future work should move beyond surface-level suppression by directly modifying internal representations—through counterfactual training, data augmentation, or representational debiasing—and by systematically linking representational divergence to behavioral outcomes, so that fairness is embedded in reasoning rather than imposed post hoc.
Extensibility. Although our evaluation focuses on natural language benchmarks, BSM readily extends to other modalities, including code generation, multilingual systems, and multimodal LLMs. We view this as a path toward a unified methodology for fairness auditing across domains, enabling systematic, reproducible comparisons that were not possible with prior scalar metrics alone. Our work also has limitations, detailed in Appendix A, including dataset scope, cost constraints, and interpretive boundaries for cross-vendor comparisons, which future research should address.
Appendix A Limitation